# ContextFormer: Redefining Efficiency in Semantic Segmentation

### v1

@@ 4.3 Experiments on ADE20K
As illustrated in Tab. 1, the proposed model is compared with the state-of-the-art approaches, consisting of efficient CNNs [25, 27, 52, 55] and lightweight ViTs [13, 32, 41, 62, 69], on the validation set of ADE20K dataset. For performance and efficiency evaluation, mIoU, GFLOPs, parameters, and latency are taken into account. The proposed model is trained for the following input resolutions: $512\times 512$ and $448\times 448$ on ADE20K dataset.
DeepLabV3+ with MobileNetV2 as encoder achieved best mIoU i.e. 38.1%. Notwithstanding, it encompasses a substantially higher computational burden measured in GFLOPs, latency contrasting with the proposed model, which achieves comparable efficacy with 25.2 less GFLOPs—representing a reduction of 97.67%—and 13.7M (88.96%) fewer parameters while just having 16ms latency, signifying a decrease in complexity.
Among the CNN-based baselines, the most popular model utilizing MobileNetV3-Large [27] as the encoder and LR-ASPP as the decoder strikes an effective balance between the computational complexity (2.0 GFLOPs, 51 ms latency) and accuracy (33.1% mIoU). Towards the effort to make this model more efficient, a reduced version, called MobileNetV3-Large-reduce by decreasing the number of feature maps of the model was used. Our proposed ContextFormer (35.0% mIoU) surpasses both the LR-ASPP models in accuracy by 1.9% and 2.7% respectively, while having much lower latency (16ms), thus proving its robustness. This represents a latency reduction of 68.63% compared to LR-ASPP and 51.52% compared to LR-ASPP reduce.
@@
ContextFormer (Ours) TPEM 34.6 0.6 1.7M
@@
In ViT-based models, HR-NAS-B [13] integrates Transformer blocks into the HRNet [59] architecture using neural architecture search, achieving an effective balance between computational complexity (2.2 GFLOPs) and accuracy (34.9% mIoU). SegFormer has shown good performance of 37.4% mIoU. Nevertheless, the proposed model achieves mIoU of 35.0% while having 92.9% and 2.1M (56.4%) fewer GFLOPs and parameters, respectively compared to SegFormer, and 0.1% higher mIoU than HR-NAS-B. Moreover, ContextFormer demonstrates a 72.7% reduction in GFLOPs compared to HR-NAS-B and a 92.9% reduction compared to SegFormer. The superior performance of ContextFormer can be attributed to the incorporation of a robust bottleneck Trans-MDC block, which synergistically leverages the strengths of both ViT and CNNs. This design facilitates the efficient extraction and representation of robust features, optimizing the trade-off between computational efficiency and model expressiveness. It is worth emphasizing that the integration of edge maps enhances performance, as evidenced by improved mIoU metric.
For further demonstrating the effectiveness of the proposed ContextFormer, we also report the results on some recent SOTA semantic segmentation approaches in different settings. It can be clearly seen in Tab. 2 that unlike other approaches, it achieves better trade-off between performance and efficiency.
@@ 4.3.2 Ablation Study
As elaborated in Tab. 3, we analyze the contribution of the Trans-MDC block to the overall performance of the model, focusing on its individual components and their collective impact. The study is divided into two parts, examining the Modulating DepthwiseConv (MDC) block’s architecture and the integration of Vision Transformer (ViT) components.
Effects of MDC Block Design: The upper half of the study highlights the incremental benefits of specific design choices within the MDC block. Starting with the baseline configuration, which includes a $3\times 3$ depth-wise convolution layer, we observe that augmenting it with an additional $1\times 1$ depth-wise convolution layer enhances the model’s segmentation accuracy by 0.2%. This improvement, though marginal, demonstrates the effectiveness of fine-grained feature extraction at multiple scales. The inclusion of a third parallel branch in the MDC block, comprising a $3\times 3$ depth-wise separable convolution, significantly boosts performance by 3.3%. This substantial improvement underscores the importance of capturing spatial hierarchies and multi-scale contextual information for robust feature representation, thus capturing the global dependencies. Additionally, the integration of a channel attention layer further refines the extracted features, albeit with a smaller but meaningful performance gain, highlighting its utility in emphasizing critical feature channels.
@@
ContextFormer (Trans-MDC) mIoU GFLOPs Parameters
@@
Effect of Trans-MDC Block and GME Map: In the lower half of the study, we evaluate the synergy between the ViT backbone and the proposed MDC block. Initially, the model demonstrates a substantial 32.7% mIoU improvement when utilizing ViT alone, even without the MDC block, validating ViT’s efficacy in enhancing semantic feature representation. However, the stepwise inclusion of depth-wise convolution layers from the MDC block markedly improves the model’s performance. The addition of each MDC component further strengthens ViT’s capability to capture local and global context, resulting in a cumulative performance improvement of approximately 2%, as detailed. These results provide compelling evidence that the MDC block not only complements ViT but also enhances its effectiveness, resulting in a robust architecture that excels in overall segmentation tasks. Furthermore, the inclusion of edge maps alongside the RGB image further amplifies the model’s performance by 0.4%, thus effectively capturing sharp and fine-grained features within the image. This proves that the edge maps act as a complementary input, guiding the segmentation process by emphasizing object boundaries and delineating intricate structures.
@@ 4.4 Results on PASCAL Context
Tab. 4 highlights the comparative performance on the PASCAL Context dataset for 59-class and 60-class configurations. DeepLabV3+ with ENet-s16 and MobileNetV3-s16 backbones achieves mIoU59/mIoU60 scores of 43.07%/39.19% and 42.34%/38.59%, respectively, but with high computational costs of 23.00 and 22.24 GFLOPs. In contrast, the proposed model delivers competitive mIoU59/mIoU60 scores of 41.85%/37.78%, with 22.51 (97.87%) fewer GFLOPs, showcasing its efficiency.
LR-ASPP achieves mIoU59/mIoU60 scores of 38.02%/35.05% using only 2.04 GFLOPs, but our model surpasses it by 3.83%/2.73% with a marginal reduction of 1.55 (75.98%) GFLOPs. Similarly, TopFormer balances performance and efficiency but is outperformed by our model, which achieves 1.46% and 1.37% higher mIoU59 and mIoU60, respectively, while reducing GFLOPs by 0.04 (7.55%). These results establish the proposed model as achieving the optimal trade-off between performance and computational efficiency.
@@ 4.5 Results on CityScapes
To further validate the efficacy of the proposed method, we summarize the results on the CityScapes dataset in Tab. 5. While PSPNet and FCN, both with MobileNetV2 encoders, achieve high accuracy, they demand 423.4 and 317.1 GFLOPs, respectively. In comparison, our model achieves 6.5% higher mIoU than FCN with a 315.9 (99.62%) GFLOPs reduction and delivers performance comparable to PSPNet with 422.2 (99.71%) fewer GFLOPs.
L-ASPP achieves the best performance among CNN-based models but at the cost of 11.4 (90.5%) additional GFLOPs. Similarly, SegFormer, a ViT-based model, achieves 71.9% mIoU but requires 17.7 GFLOPs, whereas our method matches its performance with 16.5 (93.22%) fewer GFLOPs. Additionally, our model also surpasses TopFormer with 1.5% higher mIoU at the same computational cost, demonstrating a superior balance of performance and efficiency.
@@
ContextFormer(GME) (Ours) TPEM 68.0 1.2
@@ 4.6 Results on COCO-Stuff
Tab. 6 summarizes the performance of various models on the COCO-Stuff dataset. While DeepLabV3+ with EfficientNet-s16 achieves the highest mIoU, it incurs a substantial computational cost. In contrast, the proposed model achieves comparable accuracy with 26.50 (97.8%) fewer GFLOPs, demonstrating its computational efficiency. Similarly, compared to DeepLabV3+ with MobileNetV2-s16, which consumes 25.90 GFLOPs for an mIoU of 29.88%, our model achieves similar performance with 25.30 (97.7%) fewer GFLOPs, highlighting its scalability. Additionally, our method matches PSPNet’s accuracy while significantly reducing computational demand and surpasses TopFormer with a 0.78% higher mIoU and 0.04 (6.25%) fewer GFLOPs. These results establish the proposed model as a robust and efficient solution, achieving an optimal balance between performance and computational cost for real-world applications.
@@
ContextFormer (Ours) TPEM 29.00 0.58
ContextFormer(GME) (Ours) TPEM 29.12 0.6
@@ 5 Conclusion
In this paper, we propose a hybrid architecture called ContextFormer, for the complex problem of semantic segmentation. ContextFormer takes into account the advantages of both ViT and CNN, specifically in the bottleneck part of the model. The whole pipeline is based on efficient choices. Extensive experiments on four benchmark datasets reveal that ContextFormer finds a better trade-off between performance and efficiency than the prior works. The general nature of the model makes it suitable to be added to various other methods to improve performance and efficiency. In future, the proposed method will be tested across a range of numerous other domains. Moreover, it can serve as a baseline for future research in the direction of efficient models that combine ViTs and CNNs.
@@
Mian Muhammad Naeem Abid, Nancy Mehta, Zongwei Wu, Fayaz Ali Dharejo, Radu Timofte
@@ 6 ImageNet Pre-training
For a fair comparison, we initialize the ContextFormer model using pre-trained parameters from ImageNet. As demonstrated in Fig. 4, the classification framework of ContextFormer integrates an average pooling layer followed by a linear layer, leveraging global semantic representations to generate class scores. Given the low resolution of the input images ($224\times 224$), the target resolution of the input tokens for the Trans-MDC block is configured to be $\frac{1}{64}\times\frac{1}{64}$ of the input dimensions. Quantitative results of the proposed ContextFormer model on the ImageNet-1K dataset are shown in Tab. 7.
Table 7: ContextFormer results for ImageNet classification.
@@
Figure 4: The architecture of the proposed ContextFormer model for the task of image classification.
@@ 7 Detailed Network Structure
The detailed network structure of the proposed ContextFormer model for the efficient real-time semantic segmentation task is provided in Tab. 8. Where, $dw$ and $dw.sep$ denote depth-wise and depth-wise separable convolutions. Moreover, N, H and T denote the number of blocks, number of heads and number of target channels, respectively. The input image resolution taken into account is $512\times 512$.
Table 8: Architectural details of ContextFormer model. For input resolution of $512\times 512$.
@@
Trans-MDC Block
MDC Block $3\times 3_{dw}$ 3 - 208 1
@@ 8 Details of Feed-Forward Network
In the proposed ContextFormer model, the unified output semantics from the MDC block and ViT block are passed through the Feed-Forward Network. For the Feed-Forward Network, we have integrated depth-wise convolution layer between $1\times 1$ convolution layers and to further minimize the computational complexity, expansion factor of two is incorporated which helps the Trans-MDC block to augment the capturing of global semantics. The design of Feed-Forward Network is shown in Fig. 5.
Figure 5: Design of Feed-Forward Network.
@@ 9 Visual Results
Fig. 6 shows additional visual results of the proposed ContextFormer model with original images, ground-truth, TopFormer, and ContextFormer(GME). The proposed model demonstrates superior segmentation performance on the validation set of the ADE20K benchmark dataset, highlighting its robustness.
Figure 6: Visual results on the ADE20K validation set. Our proposed model illustrates superior consistency with the ground truth segmentation results, highlighting its robustness.

### v2

@@ 4.3 Experiments on ADE20K
ContextFormer (Ours) TPEM 34.8 0.6 1.7M
@@
As illustrated in Tab. 1, the proposed model is compared with state-of-the-art approaches, consisting of efficient CNNs [25, 27, 53, 56] and lightweight ViTs [13, 31, 42, 63, 69], on the validation set of the ADE20K dataset. For performance and efficiency evaluation, mIoU, GFLOPs, parameters, and latency are taken into account. The proposed model is trained for the following input resolutions: $512\times 512$ and $448\times 448$ on the ADE20K dataset.
DeepLabV3+ with MobileNetV2 as the encoder achieved the best mIoU i.e. 38.1%. Notwithstanding, it encompasses a substantially higher computational burden measured in GFLOPs and latency, contrasting with the proposed model, which achieves comparable efficacy with 25.2 fewer GFLOPs—representing a reduction of 97.67%—and 13.7M (88.96%) fewer parameters while having just 15ms latency, signifying a decrease in complexity.
Among the CNN-based baselines, the most popular model utilizing MobileNetV3-Large [27] as the encoder and LR-ASPP as the decoder strikes an effective balance between the computational complexity (2.0 GFLOPs, 51ms latency) and accuracy (33.1% mIoU). To make this model more efficient, a reduced version, called MobileNetV3-Large-reduce, was used. Our proposed ContextFormer (35.0% mIoU) surpasses both the LR-ASPP models in accuracy by 1.9% and 2.7%, respectively, while having much lower latency (15ms), thus proving its robustness. This represents a latency reduction of 70.59% compared to LR-ASPP and 54.55% compared to LR-ASPP-reduce.
In ViT-based models, HR-NAS-B [13] integrates Transformer blocks into the HRNet [60] architecture using neural architecture search, achieving an effective balance between computational complexity (2.2 GFLOPs) and accuracy (34.9% mIoU). SegFormer has shown good performance of 37.4% mIoU. Nevertheless, the proposed model achieves an mIoU of 35.0% while having 92.9% and 2.1M (56.4%) fewer GFLOPs and parameters, respectively, compared to SegFormer, and 0.1% higher mIoU than HR-NAS-B. Moreover, ContextFormer demonstrates a 72.7% reduction in GFLOPs compared to HR-NAS-B and a 92.9% reduction compared to SegFormer. When compared to TopFormer [69], which achieves 32.8% mIoU with 0.6 GFLOPs and 1.4M parameters, the proposed model delivers a 2.2% higher mIoU while maintaining the same GFLOPs and slightly higher parameters (1.7M). Similarly, when compared to SeaFormer [59], which achieves 34.7% mIoU with 0.6 GFLOPs and 1.7M parameters, the proposed model achieves a 0.3% higher mIoU without increasing computational costs or parameters. These results highlight the superior performance of the proposed ContextFormer in balancing accuracy and efficiency.
For further demonstrating the effectiveness of the proposed ContextFormer, we also report the results on some recent SOTA semantic segmentation approaches in different settings. It can be clearly seen in Tab. 2 that unlike other approaches, our proposed model achieves a better trade-off between performance and efficiency.
@@ 4.3.2 Ablation Study
As elaborated in Tab. 3, we analyze the contribution of the Trans-BDC block to the overall performance of the model, focusing on its individual components and their collective impact. The study is divided into two parts, examining the Branched DepthwiseConv (BDC) block’s architecture and the integration of Vision Transformer (ViT) components.
Effects of BDC Block Design: The upper half of the study highlights the incremental benefits of specific design choices within the BDC block. Starting with the baseline configuration, which includes a $3\times 3$ depth-wise convolution layer, we observe that augmenting it with an additional $1\times 1$ depth-wise convolution layer enhances the model’s segmentation accuracy by 0.2%. This improvement, though marginal, demonstrates the effectiveness of fine-grained feature extraction at multiple scales. The inclusion of a third parallel branch in the BDC block, comprising a $3\times 3$ depth-wise separable convolution, significantly boosts performance by 3.3%. This substantial improvement underscores the importance of capturing spatial hierarchies and multi-scale contextual information for robust feature representation, thus capturing the global dependencies. Additionally, the integration of a channel attention layer further refines the extracted features, albeit with a smaller but meaningful performance gain, highlighting its utility in emphasizing critical feature channels.
Effect of Trans-BDC Block and GME Map: In the lower half of the study, we evaluate the synergy between the ViT backbone and the BDC block. Initially, the model demonstrates a substantial 32.7% mIoU improvement when utilizing ViT alone, even without the BDC block, validating ViT’s efficacy in enhancing semantic feature representation. However, the stepwise inclusion of depth-wise convolution layers from the BDC block markedly improves the model’s performance. The addition of each BDC component further strengthens ViT’s capability to capture local and global context, resulting in a cumulative performance improvement of approximately 2.1%, as detailed. These results provide compelling evidence that the BDC block not only complements ViT but also enhances its effectiveness, resulting in a robust architecture that excels in overall segmentation tasks.
Furthermore, the inclusion of edge maps alongside the RGB image further amplifies the model’s performance by 0.2%, thus effectively capturing sharp and fine-grained features within the image. This proves that the edge maps act as a complementary input, guiding the segmentation process by emphasizing object boundaries and delineating intricate structures. Moreover, our explicit approach can generalize to any edge detector (e.g. Canny Edge) in a zero-shot way. Tab. 4 highlights our consistent performance, with Canny introducing additional latency without providing significant improvements over Sobel, our plain approach.
@@
ContextFormer (Trans-BDC) mIoU GFLOPs Parameters
@@
Table 4: Comparison of our model including GME variants.
GME (Variations) mIoU GFLOPs Parameters Latency(ms)
Canny 31.6 0.6 1.7M 18.8
Sobel+Canny 35.0 0.6 1.7M 19.2
Sobel (Ours) 35.0 0.6 1.7M 15.1
@@ 4.4 Results on PASCAL Context
Tab. 5 highlights the comparative performance on the PASCAL Context dataset for 59-class and 60-class configurations. DeepLabV3+ with ENet-s16 and MobileNetV2-s16 backbones achieves mIoU59/mIoU60 scores of 43.07%/39.19% and 42.34%/38.59%, respectively, but with high computational costs of 23.00 and 22.24 GFLOPs. In contrast, the proposed method delivers competitive mIoU59/mIoU60 scores of 41.85%/37.78%, with 22.51 (97.87%) fewer GFLOPs, showcasing its efficiency.
LR-ASPP achieves mIoU59/mIoU60 scores of 38.02%/35.05% using only 2.04 GFLOPs, but the proposed method surpasses it by 3.83%/2.73% with a marginal reduction of 1.55 (75.98%) GFLOPs. Similarly, TopFormer balances performance and efficiency but is outperformed by the proposed method, which achieves 1.46% and 1.37% higher mIoU59 and mIoU60, respectively, while reducing GFLOPs by 0.04 (7.55%). When compared to SeaFormer, which achieves mIoU59/mIoU60 scores of 41.49%/37.27% with 0.51 GFLOPs, the proposed method achieves 0.36% and 0.51% higher mIoU59 and mIoU60, respectively, while reducing GFLOPs by 0.02 (3.92%). These results establish the proposed method as achieving the optimal trade-off between performance and computational efficiency.
@@ 4.5 Results on CityScapes
To further validate the efficacy of the proposed method, we summarize the results on the CityScapes dataset in Tab. 6. While PSPNet and FCN, both with MobileNetV2 encoders, achieve mIoU scores of 70.2 and 61.5, they require 423.4 and 317.1 GFLOPs, respectively. In contrast, the proposed method delivers 6.7% higher mIoU than FCN with a 315.9 (99.62%) GFLOPs reduction and performs comparably to PSPNet with 422.2 (99.71%) fewer GFLOPs.
L-ASPP achieves the highest mIoU of 72.7 among CNN-based models, but at the cost of 11.4 (90.48%) additional GFLOPs compared to our method. Similarly, SegFormer, a ViT-based model, achieves a comparable mIoU of 71.9 but requires 16.5 (93.22%) more GFLOPs. The proposed method also surpasses TopFormer with 1.7% higher mIoU at the same computational cost. Additionally, when compared to SeaFormer, which achieves 66.8 mIoU with 1.2 GFLOPs, the proposed method delivers 1.4% higher mIoU without additional computational cost. These results demonstrate the proposed method’s superior balance of performance and efficiency.
@@
ContextFormer(GME) (Ours) TPEM 68.2 1.2
@@ 4.6 Results on COCO-Stuff
Tab. 7 summarizes the performance of various models on the COCO-Stuff dataset. While DeepLabV3+ with EfficientNet-s16 achieves the highest mIoU of 31.45, it incurs substantial computational costs with 27.10 GFLOPs. In contrast, the proposed method achieves comparable accuracy with 26.50 (97.8%) fewer GFLOPs, demonstrating its efficiency. Similarly, compared to DeepLabV3+ with MobileNetV2-s16, which consumes 25.90 GFLOPs for an mIoU of 29.88, the proposed method achieves similar performance with 25.30 (97.7%) fewer GFLOPs, highlighting its scalability.
Additionally, the proposed method matches PSPNet’s accuracy while significantly reducing computational demand and surpasses TopFormer with a 0.92% higher mIoU and 0.04 (6.25%) fewer GFLOPs. When compared to SeaFormer, the proposed method achieves 0.02% higher mIoU with 0.02 (3.23%) fewer GFLOPs. These results establish the proposed method as a robust and efficient solution, achieving an optimal balance between performance and computational cost for real-world applications.
@@
SeaFormer [59] SeaFormer-T [59] 29.24 0.62
ContextFormer (Ours) TPEM 29.23 0.58
ContextFormer(GME) (Ours) TPEM 29.26 0.60
4.7 Object Detection
To further validate the generalization capability of our proposed method, we conducted object detection experiments using the COCO [37] dataset to evaluate its performance in a downstream task. The COCO dataset comprises 118K training images, 5K validation images, and 20K test images. For the detection framework, we employed RetinaNet [38], a one-stage model, and integrated ContextFormer(GME) as the backbone to generate a feature pyramid across multiple scales. All models were trained on the train2017 split for 12 epochs using ImageNet pretrained weights and were evaluated on the val2017 set. As evident from Tab. 8, our model excels in the detection task, showcasing the robust generalization capability of our method.
Table 8: Object detection results on COCO dataset.
Backbone mAP GFLOPs Parameters
ShuffleNetV2 [42] 25.9 161 10.4M
MobileNetV3 [27] 27.2 162 12.3M
TopFormer-T [69] 27.1 160 10.5M
SeaFormer-T [59] 31.5 160 10.9M
ContextFormer(GME) (Ours) 31.6 160 10.9M
@@ 5 Conclusion
ContextFormer is a hybrid framework that combines the strengths of CNNs and ViTs within the bottleneck to balance efficiency, accuracy, and robustness for real-time semantic segmentation. Its efficiency is driven by three synergistic modules: TPEM, the Trans-BDC block, and the FMM blocks. Extensive experiments across four benchmark datasets demonstrate that ContextFormer achieves a superior balance between accuracy and computational efficiency compared to existing methods. Its versatile design makes it adaptable to various applications, paving the way for future research into efficient ViT-CNN hybrid models across diverse domains.
@@
Mian Muhammad Naeem Abid, Nancy Mehta, Zongwei Wu, Radu Timofte
@@ 1 ImageNet Pre-training
For a fair comparison, we initialize the ContextFormer model using pre-trained parameters from ImageNet. As demonstrated in Fig. A, the classification framework of ContextFormer integrates an average pooling layer followed by a linear layer, leveraging global semantic representations to generate class scores. Given the low resolution of the input images ($224\times 224$), the target resolution of the input tokens for the Trans-BDC block is configured to be $\frac{1}{64}\times\frac{1}{64}$ of the input dimensions. Quantitative results of the proposed ContextFormer model on the ImageNet-1K dataset are shown in Tab. A.
Table A: ContextFormer results for ImageNet classification.
@@
Figure A: The architecture of the proposed ContextFormer model for the task of image classification.
@@ 2 Detailed Network Structure
The detailed network structure of the proposed ContextFormer model for the efficient real-time semantic segmentation task is provided in Tab. B. Where, $dw$ and $dw.sep$ denote depth-wise and depth-wise separable convolutions. Moreover, N, H and T denote the number of blocks, number of heads and number of target channels, respectively. The input image resolution taken into account is $512\times 512$.
Table B: Architectural details of ContextFormer model. For input resolution of $512\times 512$.
@@
Trans-BDC Block
BDC Block $3\times 3_{dw}$ 3 - 208 1
@@ 3 Details of Feed-Forward Network
In the proposed ContextFormer model, the unified output semantics from the BDC block and ViT block are passed through the Feed-Forward Network. For the Feed-Forward Network, we have integrated depth-wise convolution layer between $1\times 1$ convolution layers and to further minimize the computational complexity, expansion factor of two is incorporated which helps the Trans-BDC block to augment the capturing of global semantics. The design of Feed-Forward Network is shown in Fig. B.
Figure B: Design of Feed-Forward Network.
@@ 4 Visual Results
Fig. C shows additional visual results of the proposed ContextFormer model with original images, ground-truth, TopFormer, and ContextFormer(GME). The proposed model demonstrates superior segmentation performance on the validation set of the ADE20K benchmark dataset, highlighting its robustness.
Figure C: Visual results on the ADE20K validation set. Our proposed model illustrates superior consistency with the ground truth segmentation results, highlighting its robustness.